OpenStudy (anonymous):

A circle has a circumference of 14 ft. Its area is halved. Describe the effect on the circumference of the circle. 1)The circumference is divided by 2. 2)The circumference doubles. 3)The circumference is divided by √3 4)The circumference is divided by √2

OpenStudy (luigi0210):

@Jhannybean

OpenStudy (anonymous):

21.98? but I still dont understand.

OpenStudy (jhannybean):

One minute.

OpenStudy (jack1):

rearrange both eqns so r = r area = pi r^2 circum = 2 pi r so rearranging: assume radius of 1: C^2 = 4 * A * pi

OpenStudy (jhannybean):

i'm not understanding this . :|

OpenStudy (jack1):

so what happens to circumference if we assume first area = 10 m^2 half area = 5m^2 circumf 1 = 2 * (sqrt (10 pi)) approx 11.2 circumf 2 = 2 * (sqrt (5 pi)) approx 7.9

OpenStudy (jhannybean):

r = r?

OpenStudy (jack1):

sorry: r = r rearrange both the circum equation and the area eqn in terms of r C/(2*pi) = sqrt(A/pi)

OpenStudy (anonymous):

thanks:)

OpenStudy (jack1):

radius in terms of circumference would be the same radius from the equation in terms of area

OpenStudy (jack1):

so did you see the relationship...?

OpenStudy (jhannybean):

Nope.

OpenStudy (jhannybean):

I think i got it now.

OpenStudy (jack1):

area 2 / area 1 = 2 / 1 circumference 2 / circumference 2 = 1.41414 = sqrt 2 ... sorry, had the area ratio value backwards before

OpenStudy (jhannybean):

Your Area and Circumference have a direct relationship, so if one increases, the other does to, vise versa. You're given C = 14. and the area is reduced by half of the original area of the circle. \[\large C=2 \pi r \quad \quad \quad A=\pi r^2\]\[\large \text{Circle 1:}\]\[\large 14= 2 \pi r\]\[\large R_{1}=\frac{C}{2\pi}\]\[\large R=\frac{14}{2 \pi}\]\[\large\text{Area}=\pi r^2=\frac{14^2}{4\pi^2}* \pi= \frac{14^2}{4\cancel{\pi^2}\pi}*\cancel{\pi}=\frac{14^2}{4\pi}\]\[\large \text{now make half the area}=\frac12*\frac{14^2}{4\pi}= \frac{14^2}{8 \pi}\]\[\large\text{Circle 2}:\]\[\large \text{get new reduced r}:\]\[\large \text{Area}= \pi r^2 = \frac{14^2}{8 \pi} \implies r^2=\frac{14^2}{8 \pi} * \frac1\pi=\frac{14^2}{8\pi^2} \implies r_{2}=\frac{14}{\pi \sqrt{8}}\]\[\large \text{New Circumference}= 2 \pi r =\cancel{2 \pi}*\frac{14}{\cancel{ 2 \pi} \sqrt{2}}=\frac{14}{\sqrt{2}}\] Now we can see how the circumference changes when we reduce the area by half.

OpenStudy (jhannybean):

For circle 1 find radius find area take half the area Circle 2 : use the halved area to find NEW radius, use NEW radius to find NEW circumference.

Parth (parthkohli):

Another way is to notice that area is related to an exponent of 2. So if the perimeter/circumference increases by a factor of 2, then the area increases by a factor of 4. If p/c increases by a factor of 4, then the area increases by a factor of 16. And so we can say that if area increases/decreases by a factor of \(n\), then the p/c has decreased by a factor of \(\sqrt{n}\). Here, \(n = 2\) so the circumference is being divided by \(\sqrt2\).

Parth (parthkohli):

That is true for any regular figure.

OpenStudy (jhannybean):

And ontop of that,you can misread the question as well, it says the area is halved. This could mean that since the circumference is directly correlated with the area, a "halved" area could mean that the circumference is divided by 2. lol. 14/2 = 7 :|

OpenStudy (jhannybean):

It's a matter of saying "the area is reduced by half" to saying "the area is halved"

Parth (parthkohli):

@Jhannybean I said that the area has decreased by a factor 2.

OpenStudy (jhannybean):

No i wasn't referring to you, i was referring to the question, lol.

Parth (parthkohli):

The question isn't wrong at all.

OpenStudy (jhannybean):

 hehehe

Parth (parthkohli):

lol

OpenStudy (unklerhaukus):

\[\begin{align}\left.\begin{array}{cc}C=2\pi r&\implies&r=\frac C{2\pi}\\A=\pi r^2&\implies&r=\sqrt\frac A\pi\end{array}\right\}\implies C&=2\pi\sqrt\frac A\pi\\C'&=2\pi\sqrt\frac {A'}\pi\\A'=A/2\qquad&=2\pi\sqrt\frac {A/2}\pi\\&=\frac2{\sqrt2}\pi\sqrt\frac {A}\pi\\&=\frac C{\sqrt2}\end{align}\]

OpenStudy (jhannybean):

Rhaukus, derivatives? your way is more complicated D: or so it seems....

Parth (parthkohli):

Derivatives? I don't see a derivative anywhere.

OpenStudy (unklerhaukus):

I've used the 'primes' to denote the NEW Circumference and New Area

Parth (parthkohli):

\(C' \) is just a notation for talking about the 2nd circumference.

Parth (parthkohli):

Exactly.

OpenStudy (unklerhaukus):

What the derivative of the area with respect to radius @ParthKohli ?

Parth (parthkohli):

Its circumference lol

OpenStudy (jhannybean):

This feels like you two are trading secrets. And i'm not in on it. :(

OpenStudy (jhannybean):

Did i get it? ^_^ lol.